OpenStudy (anonymous):

What is 3 ln 3 - ln 9 expressed as a single natural logarithm?

OpenStudy (anonymous):

Hi @blacklabel can you help me?

OpenStudy (jdoe0001):

http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif <--- use rule 3 listed there then use rule 2

OpenStudy (anonymous):

Thank you @jdoe0001

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

Just remember that 3ln(3) =ln(3^3)=ln(27)

OpenStudy (anonymous):

So basically \[3 \ln 3 - \ln 9=\ ln 27-\ln 9= \ln \frac{27}{9}\]

OpenStudy (anonymous):

Thank you @BlackLabel